# Electroweak Phase Transition in Two Scalar Singlet Model with pNGB Dark Matter

### v1

@@ IV.1 Strong First Order phase transition along SM+BSM Higgs field direction
Having elucidated the general phase transition patterns for a representative set of benchmark points within this scenario, we now delve into the model’s parametric dependence on the phase transition strength. This analysis will focus on three important model parameters: $m_{H_{3}},\sin\alpha_{2}$, and $v_{\phi}$ which play a crucial role in determining the strength of the phase transition. As a representative scenario, we choose 3 benchmark points (BP1, BP2, BP3), characterized by a strong phase transition along the Standard Model (SM) Higgs direction.
@@ V Dark matter phenomenology
The various components of Eq.(48) are the following: the co-moving number density $Y_{\chi}=\frac{n_{\chi}}{s}$, where $s$ is the entropy density, $x=m_{\chi}/T$, with $T$ representing the thermal bath’s temperature. $Y^{eq}_{\chi}$ represents the equilibrium co-moving density of $\chi$. The term $\left\langle\sigma_{\left(\chi\chi\rightarrow SM,SM\right)}v\right\rangle$ refers to the thermally averaged annihilation cross section of DM into Standard Model (SM) particles. Here, $H$ denotes the Hubble parameter. The parameter $\beta$ is expressed as $\beta(T)=\frac{g_{\star}^{1/2}(T)\sqrt{g_{\rho}(T)}}{g_{s}(T)}$, with $g_{\rho}$ and $g_{s}$ representing the effective degrees of freedom (DOFs) related to the energy density and entropy density, respectively, and $g_{\star}^{1/2}=\frac{g_{s}}{\sqrt{g_{\rho}}}\left(1+\frac{1}{3}\frac{T}{g_{s}(T)}\frac{dg_{s}(T)}{dT}\right)$. Some representative Feynman diagrams in Figure 4 illustrate that the present relic abundance of the DM in the universe is mainly determined by the parameter $\lambda_{HS}$ and scalar sector mixing angles.
@@
However, BP3 and BP7 exhibit relic densities closer to the experimentally observed value. Our primary goal is to achieve a strong first-order phase transition for either or both the Standard Model (SM) and Beyond the Standard Model (BSM) Higgs fields. This requires high values of the quartic couplings $\lambda_{HS}$ and $\lambda_{S\phi}$, which increase as $v_{s}$ (the vacuum expectation value of the scalar field $s$) decreases. As shown in Table 2, higher values of $v_{s}$ generally lead to higher relic densities, approaching the observed value. Consequently, $v_{s}$ is a critical tunable parameter for achieving the desired relic density in our model. Among the BPs in Table 3 (SM case), BP3 has the highest $v_{s}$ and thus the highest relic density (0.12). Similarly, BP7 in Table 4 (BSM case) has the highest $v_{s}$ among its counterparts, with a relic density of 0.119.
@@ V.1 Direct Detection Search
Over the years, many studies on pseudo-Nambu-Goldstone boson (pNGB) dark matter have explored various theoretical frameworks and detection prospects. Investigations on direct detection mechanisms for light dark matter was studied by Essig et al. (2012) PhysRevD.85.076007 . A pNGB is an attractive candidate for dark matter as it can evade the severe limits of constraints in the recent DM direct detection experiments Abe_2021 . In general, pNGB DM of any $\mathcal{Z_{\rm 2}}$ symmetric model has vanishing direct detection cross section at the tree level in the limit $t\rightarrow 0$ Gross_2017 where $t$ is the momentum transfer of the DM with the nucleon. However, the introduction of any additional symmetry-breaking parameter will give a non-zero direct detection cross section in the limit of zero momentum transfer Kannike:2019mzk . In our model, the direct detection cross section directly depends on the $\mathcal{Z_{\rm 3}}$ breaking scale $\mu_{3}$ and also the quartic couplings $\lambda_{HS}$. If there is more than one symmetry-breaking parameter in a model, some specific choice of those parameters can give zero direct detection cross section at tree level in the limit $t\rightarrow 0$ Alanne_2020 . As our model only has one such symmetry-breaking parameter, we don’t have a cancellation mechanism of DM - nucleon interaction at the tree level for any non-zero value of the parameter $\mu_{3}$. We use the micrOMEGAs B_langer_2015 package to calculate the direct detection cross section of dark matter-nucleon interactions. The relevant Feynman diagrams for the calculation of the DD cross-section are shown in Figure 6. The expression of spin-independent direct detection differential cross section is given by Alanne_2020
@@
where the effective Higgs-nucleon coupling $f_{n}=0.3$, the nucleon mass $m_{n}=0.946$ GeV PhysRevD.85.051503 ; ALARCON2014342 ; Cline_2013 , and the effective DM-nucleon coupling $\lambda_{\rm SI}$ is given by
@@
The parameters that play a crucial role and significantly affect the relic density are the DM mass $m_{\chi}$, the vev of the complex scalar $v_{s}$, the masses of the real singlet scalars $m_{H_{2}}$, $m_{H_{3}}$ and the mixing angle $\sin{\alpha_{1}}$. Finally, we show interesting correlations among different parameters in Figure 7. The colour bar represents the variation of the mixing angle $\sin{\alpha_{1}}$. To achieve a relic density around 0.12, the choice of a high value of $v_{s}$ is preferable as mentioned earlier. Eq.(16) reveals that the parameters $\lambda_{HS}$ and $\lambda_{S\Phi}$ feature $v_{s}$ in the denominator suggesting that this quantity increases with relic density due to its inverse relationship with annihilation cross-section which makes $v_{s}$ the most adjustable parameter for achieving the desired relic density by changing its value. As mentioned earlier, direct detection cross-section depends on the $\mathcal{Z}_{3}$ breaking scale $\mu_{3}$ and the quartic coupling $\lambda_{HS}$. However, in our analysis, we express them as functions of the free parameters like $m_{\chi}$ and $v_{s}$. The spin-independent direct detection cross section $\sigma^{\rm SI}$ obtained from micrOMEGAs has been scaled by $f_{\chi}$ to compare with the upper bounds of experiments Cline_2013 :
@@ V.2 Indirect Detection Search
Various ground and space-based experiments look for the indirect signature of dark matter Ackermann_2013 ; collaboration_2015 , where they search for the signal of dark matter pair annihilation or decay in the galactic halo or outside it. Such experiments assume the existence of DM (WIMPs) in sufficient densities within certain astrophysical environments—such as the Sun, Earth, Dwarf Spheroidal Galaxies (dSph), or the Galactic Center. DM particles pair annihilate or decay in these astrophysical objects, thereby creating SM particles that undergo decay, showering, and hadronization. These long-chain processes result in the fluxes of stable SM particles such as gamma rays, neutrinos, and various light-charged particles. All these can be searched in cosmic rays as a signal of DM by ground-based telescopes such as H.E.S.S.PhysRevLett.117.111301 and CTAs Hofmann:2023fsn or satellites such as AMS PhysRevLett.113.221102 , PAMELA Adriani_2010 , and Fermi-LAT Ackermann_2015 . If DMs are produced thermally in the early universe, then currently the velocity averaged DM self-annihilation cross-section has a natural value of $\langle\sigma_{\rm ann}v\rangle\sim 3\times 10^{-26}~{}\frac{{\rm cm}^{3}}{\rm s}$. The Fermi-LAT looks for gamma-ray signals that might be produced by the annihilation of dark matter particles in dSphs, which are faint galaxies with high dark matter densities compared to their visible matter. Meanwhile, in the absence of a discernible $\gamma$ ray signal from dSphs can impose $95\%$ C.L. upper limits on the $\gamma$-ray flux from the target. Such upper limits on the $\gamma$ ray fluxes from DM annihilation can be translated to constraints in the two-dimensional plane of the DM mass and the thermally averaged DM pair-annihilation cross section $\langle\sigma v\rangle$ Ackermann_2013 .
@@ VI Gravitational Wave Spectrum
A first-order phase transition in the early universe can serve as a primordial source of gravitational wave (GW) generation. The basic characteristic of these waves are stochastic and there are three primary mechanisms for their production as detailed in Caprini et al. Caprini_2016 :
@@
Therefore, by combining these three possible sources of gravitational wave generation, the total GW energy spectrum is expressed as Ellis_2020 :
@@
A first-order phase transition is characterized by the nucleation and expansion of a spherical bubble. The interior of this bubble follows the equation of state of the broken phase. While the surrounding area remains in the symmetric phase separated by a negligibly thin bubble wall. The difference in free energy between the inside and outside of the bubble generates an effective pressure that drives the bubble’s expansion. The parameter $\alpha$ indicates the strength of the phase transition and is expressed as PhysRevD.49.2837
$\alpha=\frac{\epsilon(T_{n})}{\rho_{R}(T_{n})}.$ (55)
The limit $\alpha\to 0$ and $\alpha\to\infty$ correspond to very weak and very strong first-order phase transition respectively. One can express $\epsilon$ as Kehayias_2010
$\epsilon(T_{n})=\Delta V_{eff}(H_{i},T)-T\frac{d\Delta V_{eff}(H_{i},T)}{dT}\bigg{|}_{T=T_{n}}$ (56)
where $\Delta V_{eff}(H_{i},T)$ is the difference between effective potentials at true and false vacuums respectively and $\rho_{R}$ is the energy density of radiation which is expressed as,
$\rho_{R}(T_{n})=\frac{\pi^{2}g_{*}T_{n}^{4}}{30}$ (57)
where $g_{*}$ is relativistic degrees of freedom at $T_{n}$. The parameter $\frac{\beta}{H_{n}}$ represents the ratio of the inverse of the time taken for the phase transition to complete to the Hubble parameter value at $T_{n}$. It can be expressed as Nicolis_2004 :
@@
For $\beta/H_{n}\gg$ 1222A large $\beta/H_{n}$ allows us to neglect the Hubble expansion of the Universe during the phase-transition process., the thin wall and envelope approximation effectively describes the situation. In this model, gravitational waves are predominantly produced by the envelopes, while the overlapping regions are ignored. The contribution of bubble collisions to the GW spectrum, red-shifted up to today is then given by, Jinno_2017
@@
where the peak frequency of bubble collisions is given by, Jinno_2017
@@
Sound waves are generated as the bubble propagates and collides with each other in the plasma. The contribution of sound waves to the GW spectrum is given by, Hindmarsh_2014 ; Hindmarsh_2018 ; Hindmarsh:2017gnf ; Guo_2021
@@
The plasma is ionized and gives rise to turbulence which is magnetohydrodynamic in nature. This forms another source of gravitational waves. The contribution of magnetohydrodynamic turbulence to the GW spectrum is given by, Caprini_2009
@@
The peak frequency of magnetohydrodynamic turbulence contribution is given by, Caprini_2009
@@
Figure 10: This figure presents the gravitational wave (GW) spectrum associated with benchmark points BP1, BP4, BP5, and BP6. We emphasize the sensitivity regions of only those detectors that exhibit a detectable GW spectrum for at least one of the benchmark points. The sensitivity curves for DECIGO, U-DECIGO, and their corresponding correlation analyses are obtained from Nakayama_2010 .
@@
The generated gravitational wave would be possibly detected by the late time observatories such as LISA Bartolo_2016 , Taiji Ruan:2018tsw , TianQin Luo_2016 , Big Bang Observer (BBO)Cutler_2009 , DECi-hertz Interferometer GW Observatory (DECIGO) and Ultimate-DECIGO Kawamura_2011 . To determine the detectability of a signal against the background, the most frequently used measure is the Signal-to-Noise Ratio (SNR), defined as follows, Ellis_2020
@@
where $T$ is the total duration of the experiment. We assume $T$ to be 5 years for all relevant detectors. The quantity $h^{2}\Omega_{\text{Sens}}(f)$ represents the experimental sensitivity of a specific configuration to cosmological sources, derived from the power spectral density (PSD) $S_{h}(f)$ Kuroda_2015 .
@@
BP4 8362.47 82982 $8.51\times 10^{7}$
BP5 - 70042.9 $7.05\times 10^{7}$
BP6 - 28095.6 $2.86\times 10^{7}$

### v2

@@ IV.1 Strong First Order phase transition along SM+BSM Higgs field direction
Having elucidated the general phase transition patterns for a representative set of benchmark points within this scenario, we now delve into the model’s parametric dependence on the phase transition strength. This analysis will focus on three important model parameters: $m_{H_{3}},\sin\alpha_{2}$, and $v_{\phi}$ which play a crucial role in determining the strength of the phase transition. As a representative scenario, we choose 3 benchmark points (BP1, BP2, BP3), characterized by a strong phase transition along the SM Higgs direction.
@@ V Dark matter phenomenology
The various components of Eq.(48) are the following: the co-moving number density $Y_{\chi}=\frac{n_{\chi}}{s}$, where $s$ is the entropy density, $x=m_{\chi}/T$, with $T$ representing the thermal bath’s temperature. $Y^{eq}_{\chi}$ represents the equilibrium co-moving density of $\chi$. The term $\left\langle\sigma_{\left(\chi\chi\rightarrow SM,SM\right)}v\right\rangle$ refers to the thermally averaged annihilation cross section of DM into SM particles. Here, $H$ denotes the Hubble parameter. The parameter $\beta$ is expressed as $\beta(T)=\frac{g_{\star}^{1/2}(T)\sqrt{g_{\rho}(T)}}{g_{s}(T)}$, with $g_{\rho}$ and $g_{s}$ representing the effective degrees of freedom (DOFs) related to the energy density and entropy density, respectively, and $g_{\star}^{1/2}=\frac{g_{s}}{\sqrt{g_{\rho}}}\left(1+\frac{1}{3}\frac{T}{g_{s}(T)}\frac{dg_{s}(T)}{dT}\right)$. Some representative Feynman diagrams in Figure 4 illustrate that the present relic abundance of the DM in the universe is mainly determined by the parameter $\lambda_{HS}$ and scalar sector mixing angles.
@@
However, BP3 and BP7 exhibit relic densities closer to the experimentally observed value. Our primary goal is to achieve a strong first-order phase transition for either or both the SM and BSM Higgs fields. This requires high values of the quartic couplings $\lambda_{HS}$ and $\lambda_{S\phi}$, which increase as $v_{s}$ (the vacuum expectation value of the scalar field $s$) decreases. As shown in Table 2, higher values of $v_{s}$ generally lead to higher relic densities, approaching the observed value. Consequently, $v_{s}$ is a critical tunable parameter for achieving the desired relic density in our model. Among the BPs in Table 3 (SM case), BP3 has the highest $v_{s}$ and thus the highest relic density (0.12). Similarly, BP7 in Table 4 (BSM case) has the highest $v_{s}$ among its counterparts, with a relic density of 0.119.
@@ V.1 Direct Detection Search
Over the years, many studies on pseudo-Nambu-Goldstone boson (pNGB) dark matter have explored various theoretical frameworks and detection prospects. Investigations on direct detection mechanisms for light dark matter was studied by Essig et al. (2012) PhysRevD.85.076007 . A pNGB is an attractive candidate for dark matter as it can evade the severe limits of constraints in the recent DM direct detection experiments Abe:2021byq . In general, pNGB DM of any $\mathcal{Z_{\rm 2}}$ symmetric model has vanishing direct detection cross section at the tree level in the limit $t\rightarrow 0$ Gross:2017dan where $t$ is the momentum transfer of the DM with the nucleon. However, the introduction of any additional symmetry-breaking parameter will give a non-zero direct detection cross section in the limit of zero momentum transfer Kannike:2019mzk . In our model, the direct detection cross section directly depends on the $\mathcal{Z_{\rm 3}}$ breaking scale $\mu_{3}$ and also the quartic couplings $\lambda_{HS}$. If there is more than one symmetry-breaking parameter in a model, some specific choice of those parameters can give zero direct detection cross section at tree level in the limit $t\rightarrow 0$ Alanne:2020jwx . As our model only has one such symmetry-breaking parameter, we don’t have a cancellation mechanism of DM - nucleon interaction at the tree level for any non-zero value of the parameter $\mu_{3}$. We use the micrOMEGAs Belanger:2014vza package to calculate the direct detection cross section of dark matter-nucleon interactions. The relevant Feynman diagrams for the calculation of the DD cross-section are shown in Figure 6. The expression of spin-independent direct detection differential cross section is given by Alanne:2020jwx
@@
where the effective Higgs-nucleon coupling $f_{n}=0.3$, the nucleon mass $m_{n}=0.946$ GeV PhysRevD.85.051503 ; ALARCON2014342 ; Cline:2012hg , and the effective DM-nucleon coupling $\lambda_{\rm SI}$ is given by
@@
The parameters that play a crucial role and significantly affect the relic density are the DM mass $m_{\chi}$, the vev of the complex scalar $v_{s}$, the masses of the real singlet scalars $m_{H_{2}}$, $m_{H_{3}}$ and the mixing angle $\sin{\alpha_{1}}$. Finally, we show interesting correlations among different parameters in Figure 7. The colour bar represents the variation of the mixing angle $\sin{\alpha_{1}}$. To achieve a relic density around 0.12, the choice of a high value of $v_{s}$ is preferable as mentioned earlier. Eq.(16) reveals that the parameters $\lambda_{HS}$ and $\lambda_{S\Phi}$ feature $v_{s}$ in the denominator suggesting that this quantity increases with relic density due to its inverse relationship with annihilation cross-section which makes $v_{s}$ the most adjustable parameter for achieving the desired relic density by changing its value. As mentioned earlier, direct detection cross-section depends on the $\mathcal{Z}_{3}$ breaking scale $\mu_{3}$ and the quartic coupling $\lambda_{HS}$. However, in our analysis, we express them as functions of the free parameters like $m_{\chi}$ and $v_{s}$. The spin-independent direct detection cross section $\sigma^{\rm SI}$ obtained from micrOMEGAs has been scaled by $f_{\chi}$ to compare with the upper bounds of experiments Cline:2012hg :
@@ V.2 Indirect Detection Search
Various ground and space-based experiments look for the indirect signature of dark matter Fermi-LAT:2013thd ; Fermi-LAT:2015qzw , where they search for the signal of dark matter pair annihilation or decay in the galactic halo or outside it. Such experiments assume the existence of DM (WIMPs) in sufficient densities within certain astrophysical environments—such as the Sun, Earth, Dwarf Spheroidal Galaxies (dSph), or the Galactic Center. DM particles pair annihilate or decay in these astrophysical objects, thereby creating SM particles that undergo decay, showering, and hadronization. These long-chain processes result in the fluxes of stable SM particles such as gamma rays, neutrinos, and various light-charged particles. All these can be searched in cosmic rays as a signal of DM by ground-based telescopes such as H.E.S.S.PhysRevLett.117.111301 and CTAs Hofmann:2023fsn or satellites such as AMS PhysRevLett.113.221102 , PAMELA PAMELA:2010kea , and Fermi-LAT Fermi-LAT:2015att . If DMs are produced thermally in the early universe, then currently the velocity averaged DM self-annihilation cross-section has a natural value of $\langle\sigma_{\rm ann}v\rangle\sim 3\times 10^{-26}~{}\frac{{\rm cm}^{3}}{\rm s}$. The Fermi-LAT looks for gamma-ray signals that might be produced by the annihilation of dark matter particles in dSphs, which are faint galaxies with high dark matter densities compared to their visible matter. Meanwhile, in the absence of a discernible $\gamma$ ray signal from dSphs can impose $95\%$ C.L. upper limits on the $\gamma$-ray flux from the target. Such upper limits on the $\gamma$ ray fluxes from DM annihilation can be translated to constraints in the two-dimensional plane of the DM mass and the thermally averaged DM pair-annihilation cross section $\langle\sigma v\rangle$ Fermi-LAT:2013thd .
@@ VI Gravitational Wave Spectrum
A first-order phase transition in the early universe can serve as a primordial source of gravitational wave (GW) generation. The basic characteristic of these waves are stochastic and there are three primary mechanisms for their production as detailed in Caprini et al. MAGIC:2016xys :
@@
Therefore, by combining these three possible sources of gravitational wave generation, the total GW energy spectrum is expressed as Ellis:2020awk :
@@
A first-order phase transition is characterized by the nucleation and expansion of a spherical bubble. The interior of this bubble follows the equation of state of the broken phase. While the surrounding area remains in the symmetric phase separated by a negligibly thin bubble wall. The difference in free energy between the inside and outside of the bubble generates an effective pressure that drives the bubble’s expansion. The parameter $\alpha$ quantifies the strength of the phase transition and is defined as Giese:2020rtr ; Giese:2020znk ; Ramsey-Musolf:2024ykk :
$\alpha=\frac{{\rm D}\theta}{3\omega}$ (55)
where $\theta$ is the trace of the energy-momentum tensor of the fluid (plasma) and is given by $\rho-3p$ with $\rho$ and $p$ denoting the energy density and pressure respectively. The term ${\rm D}\theta$ is the difference of the trace in false ($\theta_{f}$) and true ($\theta_{t}$) vacuum respectively. $\omega$ is the enthalpy density in the false vacuum, and all the parameters in Eq.(55) are evaluated at the nucleation temperature $T_{n}$. For the first step of the phase transition, in the false $(f)$ and true vacuum $(t)$, the thermodynamical quantities Tian:2024ysd : pressure ($p$), energy density ($\rho$), and enthalpy density ($\omega$) are expressed as: $p_{f}=\frac{1}{3}aT^{4}$, $\rho_{f}=aT^{4}$, $\omega_{f}=\frac{4}{3}aT^{4}$, $p_{t}=-V(H_{i},T)$, $\rho_{t}=-T\frac{dV(H_{i},T)}{dT}+V(H_{i},T)$, $\omega_{t}=T\frac{dV(H_{i},T)}{dT}$ with $a=\frac{\pi^{2}g_{*}}{30}$ respectively. Putting all the relevant expressions in the Eq. (55), we get the expression of $\alpha$ in terms of the effective potential $V(H_{i},T)$ given by,
$\alpha=\frac{\frac{T}{4}\frac{dV(H_{i},T)}{dT}-V(H_{i},T)}{\frac{\pi^{2}g_{*}T^{4}}{30}}\bigg{|}_{T=T_{n}}.$ (56)
The parameter $\frac{\beta}{H_{n}}$ represents the ratio of the inverse of the time taken for the phase transition to complete to the Hubble parameter value at $T_{n}$. It can be expressed as Nicolis:2003tg :
@@
For $\beta/H_{n}\gg$ 1333A large $\beta/H_{n}$ allows us to neglect the Hubble expansion of the Universe during the phase-transition process., the thin wall and envelope approximation effectively describes the situation. In this model, gravitational waves are predominantly produced by the envelopes, while the overlapping regions are ignored. The contribution of bubble collisions to the GW spectrum, red-shifted up to today is then given by, Jinno:2016vai
@@
where the peak frequency of bubble collisions is given by, Jinno:2016vai
@@
Sound waves are generated as the bubble propagates and collides with each other in the plasma. The contribution of sound waves to the GW spectrum is given by, Hindmarsh:2013xza ; Hindmarsh:2016lnk ; Hindmarsh:2017gnf ; Guo:2020grp
@@
The plasma is ionized and gives rise to turbulence which is magnetohydrodynamic in nature. This forms another source of gravitational waves. The contribution of magnetohydrodynamic turbulence to the GW spectrum is given by, Caprini:2009yp
@@
The peak frequency of magnetohydrodynamic turbulence contribution is given by, Caprini:2009yp
@@
Figure 10: This figure presents the gravitational wave (GW) spectrum associated with benchmark points BP1, BP4, BP5, and BP6. We emphasize the sensitivity regions of only those detectors that exhibit a detectable GW spectrum for at least one of the benchmark points. The sensitivity curves for DECIGO, U-DECIGO, and their corresponding correlation analyses are obtained from Nakayama:2009ce .
@@
The generated gravitational wave would be possibly detected by the late time observatories such as LISA Bartolo:2016ami , Taiji Ruan:2018tsw , TianQin TianQin:2015yph , Big Bang Observer (BBO)Cutler:2009qv , DECi-hertz Interferometer GW Observatory (DECIGO) and Ultimate-DECIGO Kawamura:2006up . To determine the detectability of a signal against the background, the most frequently used measure is the Signal-to-Noise Ratio (SNR), defined as follows, Ellis:2020awk
@@
where $T$ is the total duration of the experiment. We assume $T$ to be 5 years for all relevant detectors. The quantity $h^{2}\Omega_{\text{Sens}}(f)$ represents the experimental sensitivity of a specific configuration to cosmological sources, derived from the power spectral density (PSD) $S_{h}(f)$ Kuroda:2015owv .
@@
BP4 - - $5.49\times 10^{6}$
BP5 - - $3.88\times 10^{6}$
BP6 - - $1.42\times 10^{6}$